# Neural Network Approach to Demand Estimation and Dynamic Pricing in Retail

### v1

@@ 2.2 Econometric Approach
This models includes the prices $p_{jt}$, the distance vector $\bar{d}_{j}$, which represents the distances from each product $j$ to all other products (including itself), and the price vector $\bar{p}_{j}$, which captures the prices of other products within the same category at time $t$. The functional form of $f$ is open to discussion; however, I adopt a polynomial form  (3) of degree 3.
@@ 5.1 Machine Learning Approach
The equation  7 is a simple linear demand function which was used in the estionation of the model. Other promising candidates are the parametrized functions such as $\mathcal{D}_{\theta_{i}}(P_{i})=\sum^{T}_{t=0}\max{(a_{t}+b_{t}P_{i},0)}$ or any combination of non-linear functions. Hence, the network on each observation estimates only two parameters. The loss function is mean squared error: $\tfrac{1}{n}\sum^{N}_{i=0}(\mathcal{D}_{\theta_{i}}(P_{i})-q_{i})^{2}$. The model receives information about the price only at the final step, where it is used to calculate the loss.

### v2

@@ 2.2 Econometric Approach
This models includes the prices $p_{jt}$, the distance vector $\bar{d}_{j}$, which represents the distances from each product $j$ to all other products (including itself), and the price vector $\bar{p}_{j}$, which captures the prices of other products within the same category at time $t$. The functional form of $f$ is open to discussion; however, I adopt a polynomial form  (3) of degree 3. The error term is normally distributed.
@@
Formally, it can be demonstrated that the variance of the price coefficient estimator is inversely related to the variance of the price of the good. Moreover, the more prices in the market are correlated (collinearity of price movements is measured in the $R^{2}$ of the regression $p$ on $f(\bar{d}_{j},\gamma)^{\top}\bar{p}_{t}$), the more the variance of the price coefficient is. It gives us theoretical grounds on the relationship between the mean squared error of the estimated coefficient and the variance of the price:
$\text{MSE}(\hat{\theta})=\mathbb{E}[(\hat{\theta}-\theta)^{2}]=\text{Var}(\hat{\theta})+\text{Bias}(\hat{\theta})^{2}$ (4)
$\text{MSE}(\hat{\theta})\propto\frac{1}{\text{Var}(p)}$
$\text{MSE}(\hat{\theta})\propto R^{2}$
Lemma. The variance of the price coefficient is inversely proportional to the variance of the price.
Proof.
The variance of the estimator $\hat{\beta}$ is $\text{Var}(\hat{\beta})=\left[(\mathbf{X}^{T}\mathbf{X})^{-1}\right]_{22}\sigma^{2}$, where $\mathbf{X}$ is the matrix of independent variables of (2).
According to Frisch–Waugh–Lovell theorem, $\hat{\beta}$ can be estimated from the regression of $q$ on $\hat{p}_{\text{res}}$, where $\hat{p}_{\text{res}}$ is the residual of the regression $p$ on other covariates. Hence, $\text{Var}(\hat{\beta})=\frac{\sigma^{2}}{\sum(\hat{p}_{\text{res}}-\bar{p}_{\text{res}})^{2}}$.
In turn, $\text{Var}(\hat{p}_{\text{res}})=(1-R^{2})\cdot\text{Var}(p)$
$\text{Var}(\hat{\beta})=\frac{\sigma^{2}}{\sum_{i=1}^{n}(p_{i}-\bar{p})^{2}\cdot(1-R^{2})}$ (5)
where $R^{2}$ refers to the regression of $p$ on other covariates. ∎
@@ 2.3 Optimization
To implement this model in practice, one should first collect sufficient historical data for training. Then, retrain the demand estimation model weekly using the accumulated data to forecast future demand. Finally, use the model’s predictions to optimize pricing by solving the corresponding optimization problem, setting prices based on forecasted demand and business objectives.
@@ 5.1 Machine Learning Approach
The equation  9 is a simple linear demand function which was used in the estionation of the model. Other promising candidates are the parametrized functions such as $\mathcal{D}_{\theta_{i}}(P_{i})=\sum^{T}_{t=0}\max{(a_{t}+b_{t}P_{i},0)}$ or any combination of non-linear functions. Hence, the network on each observation estimates only two parameters. The loss function is mean squared error: $\tfrac{1}{n}\sum^{N}_{i=0}(\mathcal{D}_{\theta_{i}}(P_{i})-q_{i})^{2}$. The model receives information about the price only at the final step, where it is used to calculate the loss.